if the root of the equation x2-ax+b=0 differ by unity



  1. a2 is equal to

  2. 1+a2 is equal to

  3. if one of the root of the equation x2-ax+b=0 be twice the other then 2a2 is equal to

  4. if the sum of the root of the equation x2-ax+b=0 is equal to the sum of the square of their reciprocal then a2 is equal to

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear jiya,

Let p ad p+1 be the roots

p+p+1=a

P(P+1)=b

p=a-1/2

(a-1)/2(a-1/2+1)=b

a^2-1=4b

a^2=4b+1

a^2+1=4b+2

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

sagarsingh24.iitd@gmail.com